The speed of a freight train is 14 km/h slower than the speed of a passenger train. The freight train travels 330 km in the same time that it takes the passenger train to travel 400 km. Find the speed of each train. Can you show me a simple way to solve this problem with numbered steps?
Answer by mananth(16946)

freight train x mph 330 miles
Passenger train (x+ 14)mph 400 miles

t=d/r 400
330 / x = / ( x+ 14 )
Cross multiply 14
330 ( x +14 ) = 400 * x
330 x + 4620 = 400 x
400 x -330 = 4620
70 x = 4620
/ 70 mph
x = 66 freight train mph
80 Passenger train mph